Missing ages. Molly’s age is twice Anita’s. If the sum of
the squares of their ages is 80, then what are their ages?

Answer by nyc_function(2741)

Molly's age = 2x
Anita's age = x
Sum of the square of their ages = (2x)^2 + x^2
Equate to 80 and solve for x.
(2x)^2 + x^2 = 80
4x^2 + x^2 = 80
5x^2 = 80
x^2 = 80/5
x^2 = 16
Let sqrt = square root for short
sqrt[x^2] = sqrt[16]
x = 4
Anita is 4 years old.
Molly is 2*4 years old or 8.